# Can Human Drivers and Connected Autonomous Vehicles Co-exist in Lane-Free Traffic? A Microscopic Simulation Perspective

### v1

@@ 4.2 Improvements using Adaptive Potential Lines Controller
Figure 10: The formation of APL corridors between HDVs for CM and FAM strategies. The CAVs following APL corridors are represented with green color. The figure represents the situation at $t=10$ min with 250 $veh/km$ and 20% HDVs. The $X_{CM}$ and $X_{AM}$ are set to 20 m in this example.
@@
Figure 11: The traffic flow rates and mean lateral speeds for different APL strategies and HDV penetrations. The vehicle density is set to 250 $veh/km$.
Figure 12: The difference of traffic flow values of APL and PL controllers under the same settings.
Figure11 compares the overall flow and the average lateral speeds of APL and PL controllers for a density of 250 $veh/km$. The first thing to note is the general trend of a significant decrease in LFT traffic flow with HDV penetration, even with the APL controller. Nonetheless, the APL controller significantly improves over the PL controller, especially in the beginning 0-50% HDVs penetration. To compare this further, Figure 12 shows the difference in traffic flow values for APL and PL controllers. For each APL strategy, the improvement generally peaks at a certain HDV penetration and then declines. The CM and NSCM provide the best improvements over PL for an HDV penetration lower than 40% with peak improvement of 2205 $veh/h$ (23.6%) achieved with NSCM for a density of 200 $veh/km$ and 30% HDV.
@@
Additionally, the improvements of NSCM are found to be more general regarding HDV penetration than CM. This means that NSCM continues to provide performance improvement for almost 10% more HDVs than CM. This is due to the way APL corridors are formed, where both longitudinally overlapping regions are first combined into a single region. Then, the lateral space available between HDVs in each region is used for APL corridors. With increasing HDV penetration, there are longitudinally more overlapping regions and less empty lateral space available for each region to form APL corridor. Consequently, since CM uses all HDVs in this process, it finds it challenging to form APL corridors and shows degraded performance earlier. This reason also applies when higher values of $X_{CM}$ are used with CM, as demonstrated in Figure 13, or when a higher vehicle density is simulated. Consequently, it is worth noting in Figure 13 that CM’s performance is improved for 400 $veh$ when lower $X_{CM}$ is used. In contrast, NSCM only considers an HDV for APL corridors when the HDV is moving slower than the surrounding vehicles. This reduces the overlapping longitudinal regions and allows NSCM to form more APL corridors than CM. This is also visible through the lateral trajectory plots in Figure 14(b), where the vehicle experiences APL corridors in NSCM much more than CM.
@@
Figure 12 also shows that the improvements of FAM and SVAM are smaller than CM and NSCM, almost half for 0-30% HDV penetration. However, FAM and SVAM are observed to be even more general than NSCM strategy in terms of HDV penetration, producing at least some improvement over PL controller upto 70% HDV penetration. This can be explained by their adaptive nature for forming APL corridors. Since they only consider the distance up to the follower CAV for the APL corridor, both strategies have fewer overlapping regions, resulting in a short but higher number of APL corridors than CM and NSCM. For smaller HDV penetration (0-30%), such short corridors would be less efficient than long corridors of CM and NSCM since many CAVs would have to laterally adjust their movements whenever they enter or exit such short APL corridors. This is visible from the higher lateral speeds of CAVs in Figure11 for FAM and SVAM as well as from the significantly larger lateral movements in the trajectory plots of Figure 14.
@@
On the other hand, these short APL corridors benefit from high HDV penetration and vehicle density. Under these conditions, the FAM and SVAM can still form APL corridors, unlike CM and NSCM strategies. This allows CAVs to overtake the blocking HDV and accelerate, making these strategies applicable to even high HDV penetration. However, as the HDV penetration goes above 70%, forming APL corridors becomes difficult for almost all APL strategies, and the performance becomes similar to the PL controller.
Figure11 also shows that the lateral movement of HDVs in the APL controller is significantly lower than in the PL controller. As described above, the APL corridors also push HDVs to form longitudinal groups via CAV movements, unlike PL where HDVs are more dispersed in between CAVs. These HDVs in APL either do not have enough speed incentives to change lateral position or do not find enough space for lateral movements among fast-moving CAVs in the APL corridors. Thus, for the 0-30% HDV penetration where the CM and NSCM form large APL corridors, the lateral movement of HDVs is minimal. However, as the number of long APL corridors decreases with increasing HDV penetration, the HDVs find more chances for lateral movements.
(a) 250 veh/km
@@
The previous section showed that the PL controller forms traffic waves with increasing HDV penetration. Therefore, it is also important to analyze the impact of the APL controller on this phenomenon. As shown in Figure 15, the APL strategy does not completely remove the traffic waves but reduces their intensity. The long APL corridors in CM and NSCM help to reduce the traffic waves the most; however, the short APL corridors in FAM and SVAM are found to be less effective against traffic waves. At a very high vehicle density (400 $veh/km$), all APL strategies struggle to avoid forming traffic waves.
@@
Overall, the results in this section show that for a relatively low number of HDVs (almost 30% HDVs), an APL strategy that favors unhindered movements of groups of CAVs like CM or NSCM produces better results. However, an APL strategy that favors the overtaking maneuvers of follower CAVs is more general and produces better results with higher HDV penetration. Nevertheless, even with the APL controller, the results indicate that penetration of at least 60% CAVs into LFT is necessary before any major benefits of LFT to the overall traffic flow start to appear.
@@ 5 Summary
This study examines the impact of HDVs on LFT using a microscopic simulation of a 1 $km$ ring road with a width of 10.2 meters. While HDVs make driving decisions based on individual benefits, CAVs use the PL controller for coordinated movements. The results indicate that a pure LFT scenario significantly improves road capacity compared to scenarios with only HDVs without lanes; a maximum flow of 20,400 $veh/h$ is observed for LFT compared to only 8,044 $veh/h$ for the latter. However, the flow improvement of LFT significantly drops as soon as HDVs are introduced into the system. Even a small penetration of HDVs, such as 5%, reduces the maximum flow by 16% (to 17,252 $veh/h$). With 10% HDVs, it drops by 32% (to 14,162 $veh/h$), and almost halves with 20% HDVs. In literature, the LFT is also marked by its characteristic feature of avoiding the formation of traffic waves. However, this study found that with the penetration of HDVs, traffic waves start to appear in LFT and worsen with higher HDV penetrations.
The study also introduced an APL controller approach to reduce the above performance drop. Unlike the simple PL controller, where the assigned PLs remain fixed, the APL controller adapts the PLs in the vicinity of HDVs. These areas with modified PLs are referred to as APL corridors. The study developed four APL variants that mainly differ on (1) the distance behind each HDV included in the APL corridor and (2) the conditions required for individual HDV whose vicinity is included in the corridor. The study found that the APL variants that favor the group movement of CAVs by forming long APL corridors perform better for a lower range of HDV penetration (up to almost 30%). In comparison, the APL strategies that favor the overtaking maneuvers of individual CAVs perform better for higher HDV penetration. Overall, the NSCM strategy for APL controller provides the best performance.
Even with the improved performance of the APL controller, the results indicate that the proportion of CAVs in LFT needs to be significantly high (almost 60%) before the practical benefits of LFT start to appear in terms of traffic flow.
@@ 6 Limitations and Future Work
Additionally, the driving behavior of individual humans varies significantly, while the study model used similar parameters for all HDVs. An important parameter for the HDVs is the reaction time and deceleration capability for calculating safe velocities. This can have a significant impact on outcomes. The same applies to CAVs, i.e., the study used a specific LFT controller and set of parameter values for the experiments. The outcomes may differ significantly if these parameters are changed. It is also possible that the impacts of HDVs are amplified due to the use of a ring road; even with a small proportion of HDVs, the CAVs repeatedly encounter the HDVs in each rotation, leading to an amplified effect on upstream vehicles.
The HDV model can be enhanced in the future to better represent human behavior in LFT scenarios. Driving simulator studies of LFT may be crucial for this purpose Sekeran et al. (2023), especially for estimating the reaction time and deceleration for HDVs in an LFT environment. The impact of HDVs should also be studied under different simulation scenarios to provide further insights into their effects. Similarly, the impact of HDVs on other LFT controllers should be investigated. It is important to note that the previously concluded penetration rate of 60% CAVs to start seeing the benefits of LFT indicates that LFT controller design should also consider boundary cases, such as handling situations when not all vehicles are CAVs. Even though APL controller reduces the impact of HDVs, there is still potential to further improve the LFT controller for these cases.

### v2

@@ 5.2 Improvements using Adaptive Potential Lines Controller
Figure 12: The formation of APL corridors between HDVs for CM and FAM strategies. The CAVs following APL corridors are represented with green color. The figure represents the situation at $t=50$ min with 200 $veh/km$ and 20% HDVs.
@@
Figure 13: The traffic flow rates and mean lateral speeds for different APL strategies and HDV penetrations. The vehicle density is set to 200 $veh/km$
(a) Constant margin based APL strategies
(b) Adaptive margin based APL strategies
Figure 14: The difference in traffic flow values of APL and PL controllers under the same settings. The shaded areas show the standard deviation.
Figure13 compares the overall flow and the average lateral speeds of APL and PL controllers for a density of 200 $veh/km$. The first thing to note is the general trend of a significant decrease in LFT traffic flow with HDV penetration, even with the APL controller. Nonetheless, the APL controller significantly improves over the PL controller, especially in the range of 5-60% HDVs penetration. To compare this further, Figure 14 shows the difference in traffic flow values for APL and PL controllers. For each APL strategy, the improvement generally peaks at a certain HDV penetration and then declines. The CM and NSCM provide the best improvements over PL for an HDV penetration lower than 40%, causing a peak average traffic flow of around 1000 $veh/h$ (8-12% improvement) with a density of 200-250 $veh/km$. It should be noted that the performance improvement depends significantly on the initial traffic situation (i.e., simulation seed) and the LFT parameter values used. For example, an additional traffic flow of 2500 $veh/h$ (24% improvement) was observed for one of the seeds with the NSCM method, 10% HDV penetration, and 200 $veh/km$.
@@
Additionally, the improvements of NSCM are found to be more general regarding HDV penetration than CM. This means that NSCM continues to provide performance improvement for a wider range of HDVs penetration than CM. This is due to the way APL corridors are formed, where both longitudinally overlapping regions are first combined into a single region. Then, the lateral space available between HDVs in each region is used for APL corridors. With increasing HDV penetration, there are longitudinally more overlapping regions and less empty lateral space available for each region to form APL corridor. Consequently, since CM uses all HDVs in this process, it finds it challenging to form APL corridors and shows degraded performance earlier. This reason also applies when higher values of $X_{CM}$ are used with CM, as demonstrated in Figure 15, or when a higher vehicle density is simulated. Consequently, it is worth noting in Figure 15 that CM’s performance is generally improved when a lower $X_{CM}$ is used. In contrast, NSCM only considers an HDV for APL corridors when the HDV is moving slower than the surrounding vehicles. This reduces the overlapping longitudinal regions and allows NSCM to form more APL corridors than CM. This is also visible through the lateral trajectory plots in Figure 16(b), where the vehicle experiences APL corridors in NSCM much more than CM.
@@
Figure 14(b) also shows that the improvements of FAM and SVAM are smaller than CM and NSCM, especially for starting 0-40% HDV penetration rate. However, FAM and SVAM are observed to be even more general than NSCM strategy in terms of HDV penetration, producing higher improvement than NSCM for scenarios with more than 50% HDVs. This can be explained by their adaptive nature for forming APL corridors. Since they only consider the distance up to the follower CAV for the APL corridor, both strategies have fewer overlapping regions, resulting in a short but higher number of APL corridors than CM and NSCM. For smaller HDV penetration (0-30%), such short corridors would be less efficient than long corridors of CM and NSCM since many CAVs would have to laterally adjust their movements whenever they enter or exit such short APL corridors. This is visible in Figure 16(a), where CAVs under FAM and SVAM experience shorter episodes of APL corridors than NSCM and CM. This also explains the higher performance deviation for starting 0-30% HDVs in Figure 14(b).
@@
On the other hand, these short APL corridors benefit from high HDV penetration and vehicle density. Under these conditions, the FAM and SVAM can still form APL corridors, unlike CM and NSCM strategies. This allows CAVs to overtake the blocking HDV and accelerate, making these strategies applicable to even high HDV penetration. This is also visible from the higher lateral speed of CAVs in Figure 13 for SVAM and FAM strategies than NSCM for HDV penetration of more than 50%.
Figure 13 also shows that the lateral movement of HDVs in the APL controller is significantly different from PL controller. With the formation of APL corridors, the HDVs find more space to change their lateral position to achieve higher speed, leading to increased mean lateral speed for HDVs under APL methods than the PL method. This is especially observed for starting 0-50% HDV penetration. If these APL corridors are sustained for longer periods, this also leads to an interesting phenomenon, as shown in Figure 12(b), where HDVs end up following each other and start to form longitudinal groups. For penetration rates higher than 50%, the proportion of HDVs experiencing more empty space due to the formation of APL corridors is smaller, and thus, the mean lateral speed of the HDVs also decreases, as shown in Figure 13.
(a) 200 veh/km
@@
The previous section also showed that the PL controller forms traffic waves with increasing HDV penetration. Therefore, it is also important to analyze the impact of the APL controller on this phenomenon. As shown in Figure 17, the APL strategy does not completely remove the traffic waves but reduces their intensity. The long APL corridors in CM and NSCM help to reduce the traffic waves the most; however, the short APL corridors in FAM and SVAM are found to be less effective against traffic waves. At a very high vehicle density (400 $veh/km$), all APL strategies struggle to avoid forming traffic waves.
@@
Overall, the results in this section show that for a relatively low number of HDVs (almost 40% HDVs), an APL strategy that favors unhindered movements of groups of CAVs like CM or NSCM produces better results. However, an APL strategy that favors the overtaking maneuvers of follower CAVs is more general and produces better results with higher HDV penetration. Nevertheless, even with the APL controller, the results indicate that penetration of at least 60% CAVs into LFT is necessary before any major benefits of LFT to the overall traffic flow start to appear.
(a) PL controller and all-CAVs scenarios.
(b) PL and NSCM controllers with 200 $veh/km$ and 20% HDVs.
(c) PL and NSCM controllers with 400 $veh/km$ and 20% HDVs
Figure 18: Comparison of acceleration and jerk values for PL and NSCM controllers for all-CAVs and scenarios 20% HDV penetration. The histogram analyzes the values of all CAVs recorded for all time steps. $j_{x}$ and $j_{y}$ represent the longitudinal and lateral jerk, respectively.
Method [CDF values in %]
Vehicle Density (veh/km) HDVs [%] TTC Threshold PL CM NSCM FAM SVAM
200 all-CAVs (0%) 1.5 s 0.001
3 s 0.007
20% 1.5 s 0.008 0.013 0.013 0.016 0.017
3 s 0.037 0.053 0.043 0.066 0.067
50% 1.5 s 0.023 0.021 0.039 0.047 0.034
3 s 0.074 0.076 0.136 0.131 0.119
400 all-CAVs (0%) 1.5 s 0.120
3 s 0.429
20% 1.5 s 0.114 0.128 0.225 0.183 0.213
3 s 0.409 0.451 0.792 0.638 0.738
50% 1.5 s 0.103 0.103 0.120 0.168 0.167
3 s 0.352 0.352 0.401 0.565 0.567
Table 3: The values of the cumulative distribution function (CDF) for time-to-collision (TTC) thresholds. The CDF values are within reasonable ranges for safety. The negative TTC values were removed before calculating CDF.
5.3 Comfort and Safety
The introduction of HDVs into the LFT environment and the APL controller increases lateral movements of CAVs. This behavior is visible in the vehicle trajectories in Figure 16. Since both comfort and safety are fundamental to traffic performance, it is necessary to evaluate whether these additional movements compromise either aspect.
Comfort is evaluated using vehicle acceleration and jerk, which should remain within accepted ranges to avoid discomfort to passengers. Figure 18 compares acceleration and jerk for different control strategies and HDV penetration rates; the NSCM method is used as a representative APL controller since it provided the best performance and also led to high-frequency lateral movements in Figure 16. In all-CAV scenarios, acceleration and jerk are smallest, with slightly wider spreads at higher vehicle densities. Introducing HDVs increases these spreads under the PL controller. For example, the standard deviation ($\sigma$) of longitudinal acceleration $\sigma(a_{x})$ rises from 0.12 $m/s^{2}$ to 0.24 $m/s^{2}$, and the longitudinal jerk $\sigma(j_{x})$ from 0.32 $m/s^{3}$ to 0.43 $m/s^{3}$ at 20% HDV penetration and 200 $veh/km$. The same behavior is observed across both axes, vehicle densities, and penetration rates. This also indicates that the smooth movement of CAVs gets hindered by HDVs, causing them to accelerate and decelerate more often, leading to increased spread of acceleration and jerk.
The NSCM method produces an even higher spread of acceleration and jerk due to more frequent formation and reconfiguration of APL corridors, forcing CAVs to adjust positions more often. For example, at 20% HDV and 400 $veh/km$, $\sigma(a_{x})$ increases from 0.25 $m/s^{2}$ to 0.37 $m/s^{2}$ and $\sigma(j_{x})$ from 0.38 $m/s^{3}$ to 0.55 $m/s^{3}$. Nonetheless, all values remain well within the hard limits used in the simulation for comfort. The $j_{x}$ may falsely appear to be uniformly distributed due to the log scale; however, $3\sigma$ limits indicate that 99% of the $j_{x}$ are well below the jerk limit of (-2,2) $m/s^{3}$.
While low jerk indicates smooth motion, it does not guarantee safe inter-vehicle spacing. Safety is therefore assessed using the time-to-collision (TTC) metric, defined as the gap between leader and follower divided by their relative speed. TTC represents the time until collision if both vehicles maintain their current speeds. Even though there are disagreements on the exact TTC thresholds for safe driving, values below 1.5 s are commonly classified as critical, and values below 3 s as cautionary Council et al. (2023). Table 3 shows the cumulative distribution function (CDF) values for TTC thresholds of 1.5 s and 3 s. In all-CAV scenarios, the proportion of leader–follower pairs with TTC less than 1.5 s is extremely low (0.001% at 200 $veh/km$ and 0.12% at 400 $veh/km$). Even with HDV penetration and different APL controllers, these percentages remain very low for both thresholds. The calculations also include the initialization period when vehicles are randomly positioned, which makes these results even more conservative.
Overall, the low TTC and jerk values indicate that the vehicles are driving at a sufficient distance from each other, and there was very limited sudden movement necessary to avoid collision. From which it can be concluded that the overall LFT setup used can potentially provide safe and comfortable driving.
@@ 6 Summary
This study examines the impact of HDVs on LFT using a microscopic simulation of a 1 $km$ ring road with a width of 10.2 meters. While HDVs make driving decisions based on individual benefits, CAVs use the PL controller for coordinated movements. The results indicate that a pure LFT scenario significantly improves road capacity compared to scenarios with only HDVs without lanes; a maximum flow of 16,700 $veh/h$ is observed for LFT compared to only 8,100 $veh/h$ for the latter. However, the flow improvement of LFT significantly drops as soon as HDVs are introduced into the system. Even a small penetration of HDVs, such as 5%, reduces the maximum flow by 20% (to 13,300 $veh/h$). The performance drop is significantly sharp for beginning 0-20% HDVs, where at 20% it drops by 40%, and becomes half at 40% HDV penetration. In literature, the LFT is also marked by its characteristic feature of avoiding the formation of traffic waves. However, this study found that with the penetration of HDVs, traffic waves start to appear in LFT and worsen with higher HDV penetrations and vehicle densities.
The study also introduced an APL controller approach to reduce the above performance drop. Unlike the simple PL controller, where the assigned PLs remain fixed, the APL controller adapts the PLs in the vicinity of HDVs. These areas with modified PLs are referred to as APL corridors. The study developed four APL variants that mainly differ on (1) the distance behind each HDV included in the APL corridor and (2) the conditions required for individual HDV whose vicinity is included in the corridor. The study found that the APL variants that favor the group movement of CAVs by forming long APL corridors perform better for a lower range of HDV penetration (up to almost 40%). In comparison, the APL strategies that favor the overtaking maneuvers of individual CAVs perform better for higher HDV penetration. Overall, the NSCM strategy for APL controller provides the best performance.
Even with the improved performance of the APL controller, the results indicate that the proportion of CAVs in LFT needs to be significantly high (almost 60%) before the practical benefits of LFT with CAVs start to appear in terms of traffic flow.
@@ 7 Limitations and Future Work
Additionally, the driving behavior of individual humans varies significantly, while the study model used similar parameters for all HDVs. The important parameters for the HDVs are the reaction time, minimum safety gap, and deceleration capability for calculating safe velocities. These can have a significant impact on outcomes. The same applies to CAVs, i.e., the study used a specific LFT controller and set of parameter values for the experiments. The outcomes may differ significantly if these parameters are changed. It is also possible that the impacts of HDVs are amplified due to the use of a ring road; even with a small proportion of HDVs, the CAVs repeatedly encounter the HDVs in each rotation, leading to an amplified effect on upstream vehicles.
The HDV model can be enhanced in the future to better represent human behavior in LFT scenarios. Driving simulator studies of LFT may be crucial for this purpose Sekeran et al. (2023), especially for estimating the reaction time and deceleration for HDVs in an LFT environment. The impact of HDVs should also be studied under different simulation scenarios to provide further insights into their effects. Similarly, the impact of HDVs on other LFT controllers should be investigated. It is important to note that the previously concluded penetration rate of 60% CAVs to start seeing the major benefits of LFT indicates that LFT controller design should also consider boundary cases, such as handling situations when not all vehicles are CAVs. Even though APL controller reduces the impact of HDVs, there is still potential to further improve the LFT controller for these cases. The inclusion of HDVs makes safety considerations even more critical, as the LFT controller cannot directly manage these vehicles. Even though the current study considered acceleration and jerk limits for improved saftey and comfort, there is significant potential to improve them further. In this context, incorporating a safe acceleration formulation for the lateral axis—similar to that used for the longitudinal axis—could be beneficial. Additionally, the future work should also consider minimizing the lateral movements of APL strategies. The frequent activation and deactivation of APL corridors can reduce passenger comfort which can be looked into in the future for forming stable and long-standing APL corridors in-between HDVs for enhanced passenger comfort.